OpenStudy (anonymous):

Gold in the form of solid cylinders will be molded into solid blocks. Each cylinder has a diameter of 6.19 cm and a height of 6 cm. Each solid block will have a volume of 360 cm3. How many cylinders of gold will it take to mold one block of gold? 1 2 3 4 5

OpenStudy (anonymous):

\[\frac{360}{\pi \left(\frac{6.19}{2}\right)^2 6}=1.99379 \]2 cylinders are required.

OpenStudy (anonymous):

Thank you robtobey :-)